If log5(log4(log3 x)) = 0 , then x = ??
Answer by nerdybill(7384)

If log5(log4(log3 x)) = 0 , then x = ??
.
Start with:
log5(log4(log3 x)) = 0
log4(log3 x) = 5^0
log4(log3 x) = 1
log3 x = 4^1
log3 x = 4
x = 3^4
x=81